# A Method to Improve the Performance of Reinforcement Learning Based on the $\mathcal{Y}$ Operator for a Class of Stochastic Differential Equation-Based Child-Mother Systems

### v3

@@ Abstract
journal: Journal of Mathematical Analysis and Applicationst1t1footnotetext: This paragraph of the first footnote will contain the date on which you submitted your paper for review.
@@ 1 INTRODUCTION
The advent of OpenAI’s Chat Generative Pre-trained Transformer (ChatGPT) marked a seminal moment in the commercialization of Large Language Models (LLMs), signaling the advent of AI’s semantic imaging prowess. This breakthrough catalyzed a paradigm shift in AI research, steering focus towards the domain of decision-making. At the forefront of this domain is reinforcement learning, which, in the wake of escalating computational capabilities, has undergone a metamorphosis into deep reinforcement learning (DRL). The seminal paper mnih2015human ignited an explosion of interest in DRL, spawning a proliferation of sophisticated algorithms such as Deep Deterministic Policy Gradient (DDPG)lillicrap2015continuous , Trust Region Policy Optimization (TRPO)schulman2015trust , Asynchronous Advantage Actor Critic (A3C)mnih2016asynchronous and Proximal Policy Optimization (PPO)schulman2017proximal . These innovations exemplify the renaissance in reinforcement learning research. DRL distinguishes itself by utilizing deep neural networks to distill features from systemic statesbohmer2015autonomous , thus enabling a seamless bridge between action and state spacestan2019energy and facilitating end-to-end system optimization. This method addresses the limitations inherent in traditional control mechanismssong2023reaching and has seen extensive exploration and implementation in roboticssalvato2021crossing and autonomous vehicular controlli2022decision . Despite the advancements in reinforcement learning (RL), its application remains constrained by the exigencies of environmental modeling during the training phase. Both model-based and model-free algorithms presuppose a level of environmental determinism that belies the inherent randomness of real-world settings. This discordance reveals that proficiency within simulated environments does not necessarily translate to real-world efficacy, as highlighted in canese2021multi . Addressing this disjunction, researchers are exploring avenues to obviate the need for precise environmental modeling. One school of thought, as elucidated in liu2021policy , advocates for direct agent-environment interactions, either within the actual milieu or subsequent to simulated training. However, this presents considerable safety risks accentuated by the opaque nature of the neural networks governing end-to-end deep RL, which is in stark contrast to the mathematical transparency of traditional control methods. This concern has propelled research into safety reinforcement learning, with significant contributions by thananjeyan2021recovery , stooke2020responsive and cheng2019end marvi2021safe . An alternative approach posits the integration of environmental uncertainty into simulations, with the intention of fostering more adaptable agents. Techniques such as Gaussian processesengel2005reinforcement and Stochastic Differential Equations (SDEs)yang2023parameters are employed to model this uncertainty. Recent innovations leverage deep neural networks to ascertain the drift and diffusion terms of SDEs from extensive environmental dataxu2022infinitely yang2023neural . This approach called neural stochastic differential equations(NSDEs) that has gained traction in diverse fields, from financechen2021deep to unmanned aircraft systemsdjeumou_how_2023 and autonomous drivingqi_stochastic_2022 . Furthermore, the confluence of SDE-based models and RL methodologies is yielding novel solutions to optimal control challengeswang2020reinforcement wang2020continuous wang1812exploration . The zhang2022online advanced a stochastic model for n-player non-zero sum differential games, demonstrating convergence to Nash equilibria via Q-learning. The pirmorad2021deep employed RL to modulate a 1-dimensional Stochastic Burgers’ equation, successfully dampening shock waves and mitigating sharp gradients. Moreover, chen_incremental_2019 introduced an AC based RL framework for SDE-modeled systems, proving the convergence of the Critic network and proposing a simplified surrogate function to streamline computations. SDEs have been adeptly employed to capture the inherent stochasticity of controlled systems, presenting a distinct advantage over traditional models predicated on Ordinary Differential Equations (ODEs). This approach, showcased in zhang2022online pirmorad2021deep chen_incremental_2019 , offers a more nuanced representation of real-world variability. Despite these advancements, two pertinent issues emerge: Problem I: The integration of stochasticity in SDE-based models does not extend to the design of the state-value function estimation in AC based reinforcement learning architectures. Traditional deep RL algorithms design the Critic network’s loss function to minimize the Mean Square Error (MSE) between the estimated value function at the current and subsequent time steps, as represented by the equations:
@@
In studies such as those by zhang2022online and pirmorad2021deep , stochasticity is confined to the controlled system’s model, with the design of RL algorithms—both Actor and Critic networks—failing to accommodate this variability. Consequently, the optimization of policy relies heavily on the generalization ability of RL methods. Moreover, chen_incremental_2019 simplifies the Critic network’s loss function, derived from the Itô diffusion process’s characteristic operator, to a surrogate function similar to the traditional form shown in (1) for computational efficiency, thereby overlooking the system’s stochastic aspects. wang2020continuous also primarily concentrate on the Actor network’s design while employing SDEs to model the system.
@@
Problem II: The utilization of SDEs in the modeling of control systems, as evidenced in works like chen_incremental_2019 , imposes stringent continuity conditions on the state-value function—specifically, the necessity for at least second-order continuous and $\alpha$-$\ddot{\text{o}}$lder continuity. This prerequisite stems from the optimization methodologies, such as the Hamilton-Jacobi-Bellman (HJB) equation or reinforcement learning algorithms, which require the state-value function’s partial derivatives for optimal policy determination. This requirement restricts the effectiveness of these methods in two distinct cases:
@@
Addressing these challenges, this paper introduces an alternative operator, referred to as the $\mathcal{Y}$ operator, that is functionally equivalent to the characteristic operator of the Itô diffusion process. A rigorous proof leveraging the Kolmogorov forward equation and Gaussian distribution properties is provided in the 3 section. Utilizing the $\mathcal{Y}$ operator within the AC framework, we design a reinforcement learning controller for the child-mother system detailed in the 4 section. The $\mathcal{Y}$ operator not only resolves the issue of disregarding the controlled system’s stochasticity in value function estimation, as stated in Problem I, but also circumvents the need for computing the value function’s partial derivatives when deriving optimal policy, as highlighted in Problem II. This is achieved by $\mathcal{Y}$ operator transforming the problem of calculating the value function’s partial derivatives into one of computing the drift and diffusion terms’ partial derivatives in the system’s describing SDEs. It provides insightful solutions for the special tasks mentioned above, including those requiring diverse activation function choices in neural network design, inverse reinforcement learning tasks, and problems agnostic to reward functions. The child-mother system, widely applicable in autonomous driving, is formulated as follows:
@@
where $z_{t}$ is the child-system state, $z_{t}$ is affected only by the child-system state $z_{t}$ and the control input $u_{t}$, and $w_{t}$ is the mother-system state, $w_{t}$ is affected only by the child-system state $z_{t}$ and the mother-system state $w_{t}$. This type of system has a wide prospect in the field of autonomous driving, where the child-system can be considered as the ego vehicle that can be controlled and the mother-system as all surrounding vehicles excluding the ego vehicle. Most of the existing studies in the field of autonomous driving divide the child-mother system into two systemsgu2022integrated , the child-system (i.e., the ego vehicle in autonomous driving) is modeled by a two-wheeled bicycle model and a non-linear tire modelsnider2009automatic , and the mother-system (i.e., the surrounding vehicles in autonomous driving) is modeled by the Intelligent Driver Model(IDE)treiber2000congested , the Minimizing Overall Braking Induced by Lane change(MOBIL) modeltreiber2016mobil , etc. However, this type of approach does not reasonably consider the stochasticity in the real model, and in the 4 section of this article, this child-mother system is modeled using data-driven NSDEs, which fully considers the uncertainty in the real environment and also gives how to calibrate the deep neural network parameters of this child-mother system using a large amount of data collected in a realistic environment. This paper proposes a novel reinforcement learning framework called $\mathcal{Y}$ Operator-based Reinforcement Learning (YORL). It contrasts this with the Traditional Stochastic Reinforcement Learning (TSRL) methodology, where the Critic network’s loss function is conventionally designed. YORL employs the $\mathcal{Y}$ operator to formulate a novel Critic network loss function for the child-mother system, as explicated in the 5 section. The 6 section compares the YORL and TSRL methods’ performance on this this class of child-mother systems. In summary, this paper consists of the following sections. 2 contains some theoretical knowledge in SDEs, which will be used in the subsequent deductions and proofs. 3 contains the theoretical part, which proposes the $\mathcal{Y}$ operator that is equivalent to the characteristic operator of the Itô diffusion process, gives a rigorous proof of the equivalence of these two operators by using the Kolmogorov forward equation and the properties of the Gaussian distribution, and finally gives two propositions of the $\mathcal{Y}$ operator. The proof of these propositions will be shown in the Appendix. The system studied in this paper, a special class of child-mother systems, is described in 4. 4 details the modeling of this child-mother system using SDEs, and gives a extremely trivial method to calibrate this stochastic model using data. In 5, the $\mathcal{Y}$ operator proposed in 3 is applied to the design of the loss function of the Critic network for reinforcement learning, and build a noval reinforcement learning framework based on $\mathcal{Y}$ operator(i.e., YORL). In 6, the reinforcement learning framework designed in 5 is simulated in linear and nonlinear child-mother system, respectively, to solve the optimal control problem and compared with the TSRL method. 7 summarize the full paper and provide some outlooks for the future.
@@ 2 PRELIMINARIES
{evans2012introduction Section4.2: The future information} Let $W(\cdot)$ be a $m$-dimensional Brownian motion defined on some probability space $(\Omega,\mathcal{F},P)$. The $\sigma$-algebra
@@
{evans2012introduction Section4.2: Filtration} In a probability space $(\Omega,\mathcal{F},P)$, a family $\{\mathcal{F}_{t},t\geq 0\}$ of $\sigma$-algebras have $\forall t\geq 0,\mathcal{F}_{t}\subset\mathcal{F}$. $\{\mathcal{F}_{t},t\geq 0\}$ is called filtration if there have $\forall 0\leq t\leq s,\mathcal{F}_{t}\subset\mathcal{F}_{s}$
@@
{evans2012introduction Section4.2: Progressive process} In a probability space $(\Omega,\mathcal{F},P)$, there have a filtration $\mathcal{F}_{t}$ on this probability space. The mapping $G:\Omega\times[0,T]\rightarrow\mathbb{R}^{n}$ is a progressive process iff $G(\omega,t)$ is measurable in space $\mathcal{F}_{t}\times\mathcal{B}([0,T])$, $\forall t\in[0,T]$.
@@
{evans2012introduction Section4.2: Two special spaces} (i) $\mathbb{L}^{2}_{n}(0,T)$ is the space of all real-valued, progressive processes $G:\Omega\times[0,T]\rightarrow\mathbb{R}^{n}$ such that
@@
Definition 1$\sim$4 will be used in Lemma 1 and are fundamentals to the subsequent description of SDEs.
Lemma 1.
{evans2012introduction section5.2: Existence and uniqueness of solution of stochastic differential equation} Support that $h:\mathbb{R}^{n}\times[0,T]\rightarrow\mathbb{R}^{n}$ and $H:\mathbb{R}^{n}\times[0,T]\rightarrow\mathbb{R}^{n\times m}$ are continuous and satisfy the following conditions:
$(i)\begin{cases}&\forall t\in[0,T],x,\hat{x}\in\mathbb{R}^{n},\exists L\in\mathbb{R}\text{ such that}\\ \qquad&\lvert h(x,t)-h(\hat{x},t)\rvert\leq L\lvert x-\hat{x}\rvert,\\ &\lvert H(x,t)-H(\hat{x},t)\rvert\leq L\lvert x-\hat{x}\rvert,\\ \end{cases}$ (6)
$(ii)\begin{cases}&\forall t\in[0,T],x\in\mathbb{R}^{n},\exists K\in\mathbb{R}\text{ such that}\\ \qquad&\lvert h(x,t)\rvert\leq K(1+\lvert x\rvert),\\ &\lvert H(x,t)\rvert\leq K(1+\lvert x\rvert),\\ \end{cases}$ (7)
@@
where $W(\cdot)$ is a given $m$-dimensional Brownian motion. Then there exists a unique solution $X\in\mathbb{L}^{2}_{n}(0,T)$ of the SDE:
@@ 3 THEORY
{evans2012introduction section4.4: Itôz’s chain rule} Considering a mapping $\Psi:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a continuous, with continuous partial derivatives $\frac{\partial\Psi(X_{t})}{\partial X_{t,i}}$ and $\frac{\partial^{2}\Psi(X_{t})}{\partial x_{i}\partial x_{j}}$ mapping, for $i,j=1,2,\dots,n$. The Itô chain rule is
@@
where $(\cdot)_{i}$, $(\cdot)_{i,j}$ are the element of row $i$ of the matrix and the element of column $j$ of row $i$ of the matrix, respectively. Meanwhile, the expectation of (12) can be written as
@@
{oksendal2003stochastic section7.5: Characteristic Operators of Itô Diffusion Process} Considering Lemma 2, introduce a characteristic operator $\mathcal{A}_{X}$ for the Itô diffusion process. For a stohastic process $\{X_{t}\}$ in (11) and a mapping $\Psi$ in (12), characteristic operator $\mathcal{A}_{X}$ can be defined as
@@
{oksendal2003stochastic section7.3: The generator of Itô Diffusion} Assuming a random variable $B$, the probability density function of it is denoted as $p(b)$. Denoted
[…]
@@
{ludvigsson2013kolmogorov section3.2: Kolmogorov Forward Equation} According to Kolmogorov forward equationludvigsson2013kolmogorov , the characteristic operator $\mathcal{A}_{X}$ and the dual form $\mathcal{A}^{*}_{X}$ satisfy the following relationship
@@
($\mathcal{Y}$ operator) The $\mathcal{Y}$ operator with respect to the stochastic process $\{X_{t}\}$ can be formulated as
@@
where $X_{t-}$ is the state of the stochastic process $\{X_{t}\}$ at an extremely short moment before time $t$, i.e., the history information of the last extremely short moment of $X_{t}$. Moreover, $\varLambda$ is the all-one vector and the dimension of it is equal to the dimension of $X_{t}$ which is $n$. The operator $\odot$ is the Hadamard product.
@@
(Operator Equivalence Theorem) When the $\mathcal{Y}$ operator is given well-defined as described in Definition 6, the $\mathcal{Y}$ operator with respect to the stochastic process $\{X_{t}\}$, which is $\mathcal{Y}_{X}$, is equivalent to the characteristic operator of Itô diffusion process $\mathcal{A}_{X}$.
@@
According to Corollary 2, $<\mathcal{A}_{X}\Psi(X_{t}),p(X_{t})>$ can be formulated as
@@
where $I_{E}$ is the identity matrix. Therefore, $<\mathcal{A}_{X}\Psi(X_{t}),p(X_{t})>$ can be rewritten as
@@
Assuming $X^{(1)}_{t}$, $X^{(2)}_{t}$dots $X^{(m)}_{t}$ are obey Itô diffusion process as (11)
@@
If $dW^{(1)}_{t}$, $dW^{(2)}_{t}$dots $dW^{(m)}_{t}$ are independent of each other and there exists a function $\Psi(X^{(1)}_{t},X^{(2)}_{t}\dots X^{(m)}_{t})$ of $X^{(1)}_{t}$, $X^{(2)}_{t}$dots $X^{(m)}_{t}$. Then the expectation of the derivative with respect to time of a multivariate continuous function $\Psi(X^{(1)}_{t},X^{(2)}_{t}\dots X^{(m)}_{t})$ satisfying the conditions in Lemma 2 is of the form
[…]
@@
Proofs of Propositions 1, 3 are shown in the Appendix section of the article By Corollary 2 and Proposition 3, it is clear that the problem of solving the derivative of the function $\Psi(X_{t})$ with respect to time $t$ is transformed, by two transformations of the $\mathcal{A}$ operator and the $\mathcal{Y}$ operator, into a problem of solving the partial differential equations for the drift and diffusion term functions of the SDEs obeyed by the independent variables in the function $\Psi(X_{t})$ with respect to their independent variables which means that solve for the three partial differential matrices $\mathcal{C}^{1}_{h,x,t}$, $\mathcal{C}^{1}_{H,x,t}$ and $\mathcal{C}^{2}_{H,x,t}$ in the $\mathcal{Y}$ operator. Furthermore, the transformation from the $\mathcal{A}$ operator to the $\mathcal{Y}$ operator is essentially a transformation from a partial differential operator of the function $\Psi(X_{t})$ to a linear operator of the function $\Psi(X_{t})$. the $\mathcal{Y}$ operator, as a linear operator of the function $\Psi(X_{t})$, would have much better properties, both in mathematics and engineering.

### v4

@@ Abstract
journal: Mathematics and Computers in Simulationt1t1footnotetext: This paragraph of the first footnote will contain the date on which you submitted your paper for review.
@@ 1 INTRODUCTION
The advent of OpenAI’s Chat Generative Pre-trained Transformer (ChatGPT) marked a seminal moment in the commercialization of Large Language Models (LLMs), signaling the advent of AI’s semantic imaging prowess. This breakthrough catalyzed a paradigm shift in AI research, steering focus towards the domain of decision-making. At the forefront of this domain is reinforcement learning, which, in the wake of escalating computational capabilities, has undergone a metamorphosis into deep reinforcement learning (DRL). The seminal paper [1] ignited an explosion of interest in DRL, spawning a proliferation of sophisticated algorithms such as Deep Deterministic Policy Gradient (DDPG)[2], Trust Region Policy Optimization (TRPO)[3], Asynchronous Advantage Actor Critic (A3C)[4] and Proximal Policy Optimization (PPO)[5]. These innovations exemplify the renaissance in reinforcement learning research.
DRL distinguishes itself by utilizing deep neural networks to distill features from systemic states[6] , thus enabling a seamless bridge between action and state spaces[7] and facilitating end-to-end system optimization. This method addresses the limitations inherent in traditional control mechanisms[8] and has seen extensive exploration and implementation in robotics[9] and autonomous vehicular control[10].
Despite the advancements in reinforcement learning (RL), its application remains constrained by the exigencies of environmental modeling during the training phase. Both model-based and model-free algorithms presuppose a level of environmental determinism that belies the inherent randomness of real-world settings. This discordance reveals that proficiency within simulated environments does not necessarily translate to real-world efficacy, as highlighted in [11]. Addressing this disjunction, researchers are exploring avenues to obviate the need for precise environmental modeling. One school of thought, as elucidated in [12], advocates for direct agent-environment interactions, either within the actual milieu or subsequent to simulated training. However, this presents considerable safety risks accentuated by the opaque nature of the neural networks governing end-to-end deep RL, which is in stark contrast to the mathematical transparency of traditional control methods. This concern has propelled research into safety reinforcement learning, with significant contributions by [13], [14] and [15][16].
An alternative approach posits the integration of environmental uncertainty into simulations, with the intention of fostering more adaptable agents. Techniques such as Gaussian processes[17] and Stochastic Differential Equations (SDEs)[18] are employed to model this uncertainty. Recent innovations leverage deep neural networks to ascertain the drift and diffusion terms of SDEs from extensive environmental data[19][20]. This approach called neural stochastic differential equations(NSDEs) that has gained traction in diverse fields, from finance[21] to unmanned aircraft systems[22] and autonomous driving[23].
Furthermore, the confluence of SDE-based models and RL methodologies is yielding novel solutions to optimal control challenges[24][25][26]. The [27] advanced a stochastic model for n-player non-zero sum differential games, demonstrating convergence to Nash equilibria via Q-learning. The [28] employed RL to modulate a 1-dimensional Stochastic Burgers’ equation, successfully dampening shock waves and mitigating sharp gradients. Moreover, [29] introduced an AC based RL framework for SDE-modeled systems, proving the convergence of the Critic network and proposing a simplified surrogate function to streamline computations.
SDEs have been adeptly employed to capture the inherent stochasticity of controlled systems, presenting a distinct advantage over traditional models predicated on Ordinary Differential Equations (ODEs). This approach, showcased in [27][28][29], offers a more nuanced representation of real-world variability. Despite these advancements, two pertinent issues emerge:
Problem I: The integration of stochasticity in SDE-based models does not extend to the design of the state-value function estimation in AC based reinforcement learning architectures.
Traditional deep RL algorithms design the Critic network’s loss function to minimize the Mean Square Error (MSE) between the estimated value function at the current and subsequent time steps, as represented by the equations:
@@
In studies such as those by [27] and [28], stochasticity is confined to the controlled system’s model, with the design of RL algorithms—both Actor and Critic networks—failing to accommodate this variability. Consequently, the optimization of policy relies heavily on the generalization ability of RL methods. Moreover, [29] simplifies the Critic network’s loss function, derived from the Itô diffusion process’s characteristic operator, to a surrogate function similar to the traditional form shown in (1) for computational efficiency, thereby overlooking the system’s stochastic aspects. [25] also primarily concentrate on the Actor network’s design while employing SDEs to model the system.
@@
[…]
@@
Addressing these challenges, this paper introduces an alternative operator, referred to as the $\mathcal{Y}$ operator, that is functionally equivalent to the characteristic operator of the Itô diffusion process. A rigorous proof leveraging the Kolmogorov forward equation and Gaussian distribution properties is provided in the 3 section. Utilizing the $\mathcal{Y}$ operator within the AC framework, we design a reinforcement learning controller for the child-mother system detailed in the 4 section.
The $\mathcal{Y}$ operator not only resolves the issue of disregarding the controlled system’s stochasticity in value function estimation, as stated in Problem I, but also circumvents the need for computing the value function’s partial derivatives when deriving optimal policy, as highlighted in Problem II. This is achieved by $\mathcal{Y}$ operator transforming the problem of calculating the value function’s partial derivatives into one of computing the drift and diffusion terms’ partial derivatives in the system’s describing SDEs. It provides insightful solutions for the special tasks mentioned above, including those requiring diverse activation function choices in neural network design, inverse reinforcement learning tasks, and problems agnostic to reward functions.
The child-mother system, widely applicable in autonomous driving, is formulated as follows:
@@
where $z_{t}$ is the child-system state, $z_{t}$ is affected only by the child-system state $z_{t}$ and the control input $u_{t}$, and $w_{t}$ is the mother-system state, $w_{t}$ is affected only by the child-system state $z_{t}$ and the mother-system state $w_{t}$. This type of system has a wide prospect in the field of autonomous driving, where the child-system can be considered as the ego vehicle that can be controlled and the mother-system as all surrounding vehicles excluding the ego vehicle. Most of the existing studies in the field of autonomous driving divide the child-mother system into two systems[30], the child-system (i.e., the ego vehicle in autonomous driving) is modeled by a two-wheeled bicycle model and a non-linear tire model[31], and the mother-system (i.e., the surrounding vehicles in autonomous driving) is modeled by the Intelligent Driver Model(IDE)[32], the Minimizing Overall Braking Induced by Lane change(MOBIL) model[33], etc. However, this type of approach does not reasonably consider the stochasticity in the real model, and in the 4 section of this article, this child-mother system is modeled using data-driven NSDEs, which fully considers the uncertainty in the real environment and also gives how to calibrate the deep neural network parameters of this child-mother system using a large amount of data collected in a realistic environment.
This paper proposes a novel reinforcement learning framework called $\mathcal{Y}$ Operator-based Reinforcement Learning (YORL). It contrasts this with the Traditional Stochastic Reinforcement Learning (TSRL) methodology, where the Critic network’s loss function is conventionally designed. YORL employs the $\mathcal{Y}$ operator to formulate a novel Critic network loss function for the child-mother system, as explicated in the 5 section. The 6 section compares the YORL and TSRL methods’ performance on this this class of child-mother systems.
In summary, this paper consists of the following sections. 2 contains some theoretical knowledge in SDEs, which will be used in the subsequent deductions and proofs. 3 contains the theoretical part, which proposes the $\mathcal{Y}$ operator that is equivalent to the characteristic operator of the Itô diffusion process, gives a rigorous proof of the equivalence of these two operators by using the Kolmogorov forward equation and the properties of the Gaussian distribution, and finally gives two propositions of the $\mathcal{Y}$ operator. The proof of these propositions will be shown in the Appendix. The system studied in this paper, a special class of child-mother systems, is described in 4. 4 details the modeling of this child-mother system using SDEs, and gives a extremely trivial method to calibrate this stochastic model using data. In 5, the $\mathcal{Y}$ operator proposed in 3 is applied to the design of the loss function of the Critic network for reinforcement learning, and build a noval reinforcement learning framework based on $\mathcal{Y}$ operator(i.e., YORL). In 6, the reinforcement learning framework designed in 5 is simulated in linear and nonlinear child-mother system, respectively, to solve the optimal control problem and compared with the TSRL method. 7 summarize the full paper and provide some outlooks for the future.
@@ 2 PRELIMINARIES
{[34]Section4.2: The future information}
Let $W(\cdot)$ be a $m$-dimensional Brownian motion defined on some probability space $(\Omega,\mathcal{F},P)$. The $\sigma$-algebra
@@
{[34]Section4.2: Filtration}
In a probability space $(\Omega,\mathcal{F},P)$, a family $\{\mathcal{F}_{t},t\geq 0\}$ of $\sigma$-algebras have $\forall t\geq 0,\mathcal{F}_{t}\subset\mathcal{F}$. $\{\mathcal{F}_{t},t\geq 0\}$ is called filtration if there have $\forall 0\leq t\leq s,\mathcal{F}_{t}\subset\mathcal{F}_{s}$
@@
{[34]Section4.2: Progressive process}
In a probability space $(\Omega,\mathcal{F},P)$, there have a filtration $\mathcal{F}_{t}$ on this probability space. The mapping $G:\Omega\times[0,T]\rightarrow\mathbb{R}^{n}$ is a progressive process iff $G(\omega,t)$ is measurable in space $\mathcal{F}_{t}\times\mathcal{B}([0,T])$, $\forall t\in[0,T]$.
@@
{[34]Section4.2: Two special spaces}
(i) $\mathbb{L}^{2}_{n}(0,T)$ is the space of all real-valued, progressive processes $G:\Omega\times[0,T]\rightarrow\mathbb{R}^{n}$ such that
$E\left(\int_{0}^{T}G^{2}dt\right)<\infty.$ (4)
$E\left(\int_{0}^{T}\lvert F\rvert dt\right)<\infty.$ (5)
@@
Lemma 1.
{[34]section5.2: Existence and uniqueness of solution of stochastic differential equation}
Support that $h:\mathbb{R}^{n}\times[0,T]\rightarrow\mathbb{R}^{n}$ and $H:\mathbb{R}^{n}\times[0,T]\rightarrow\mathbb{R}^{n\times m}$ are continuous and satisfy the following conditions:
@@
[…]
@@
where $W(\cdot)$ is a given $m$-dimensional Brownian motion.
Then there exists a unique solution $X\in\mathbb{L}^{2}_{n}(0,T)$ of the SDE:
@@ 3 THEORY
{[34]section4.4: Itôz’s chain rule}
Considering a mapping $\Psi:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a continuous, with continuous partial derivatives $\frac{\partial\Psi(X_{t})}{\partial X_{t,i}}$ and $\frac{\partial^{2}\Psi(X_{t})}{\partial x_{i}\partial x_{j}}$ mapping, for $i,j=1,2,\dots,n$. The Itô chain rule is
@@
where $(\cdot)_{i}$, $(\cdot)_{i,j}$ are the element of row $i$ of the matrix and the element of column $j$ of row $i$ of the matrix, respectively.
Meanwhile, the expectation of (12) can be written as
@@
{[35]section7.5: Characteristic Operators of Itô Diffusion Process}
Considering Lemma 2, introduce a characteristic operator $\mathcal{A}_{X}$ for the Itô diffusion process. For a stohastic process $\{X_{t}\}$ in (11) and a mapping $\Psi$ in (12), characteristic operator $\mathcal{A}_{X}$ can be defined as
@@
{[35]section7.3: The generator of Itô Diffusion}
Assuming a random variable $B$, the probability density function of it is denoted as $p(b)$. Denoted
$B,p(b)>=\mathbb{E}\{B\}=\int_{\mathbb{R}}Bp(b)db$ (17)
@@
{[36]section3.2: Kolmogorov Forward Equation}
According to Kolmogorov forward equation[36], the characteristic operator $\mathcal{A}_{X}$ and the dual form $\mathcal{A}^{*}_{X}$ satisfy the following relationship
$\mathcal{A}_{X}\Psi(X_{t}),p(X_{t})>=<\Psi(X_{t}),\mathcal{A}^{*}_{X}p(X_{t})>$ (19)
@@
($\mathcal{Y}$ operator)
The $\mathcal{Y}$ operator with respect to the stochastic process $\{X_{t}\}$ can be formulated as
@@
where $X_{t-}$ is the state of the stochastic process $\{X_{t}\}$ at an extremely short moment before time $t$, i.e., the history information of the last extremely short moment of $X_{t}$.
Moreover, $\varLambda$ is the all-one vector and the dimension of it is equal to the dimension of $X_{t}$ which is $n$. The operator $\odot$ is the Hadamard product.
@@
(Operator Equivalence Theorem)
When the $\mathcal{Y}$ operator is given well-defined as described in Definition 6, the $\mathcal{Y}$ operator with respect to the stochastic process $\{X_{t}\}$, which is $\mathcal{Y}_{X}$, is equivalent to the characteristic operator of Itô diffusion process $\mathcal{A}_{X}$.
$\mathcal{A}_{X}\Psi(X_{t}),p(X_{t})>=<\mathcal{Y}_{X}\Psi(X_{t}),p(X_{t})>$ (23)
@@
According to Corollary 2, $\mathcal{A}_{X}\Psi(X_{t}),p(X_{t})>$ can be formulated as
@@
where $I_{E}$ is the identity matrix.
Therefore, $\mathcal{A}_{X}\Psi(X_{t}),p(X_{t})>$ can be rewritten as
@@
Assuming $X^{(1)}_{t}$, $X^{(2)}_{t} \dots$ $X^{(m)}_{t}$ are obey Itô diffusion process as (11)
@@
If $dW^{(1)}_{t}$, $dW^{(2)}_{t} \dots$ $dW^{(m)}_{t}$ are independent of each other and there exists a function $\Psi(X^{(1)}_{t},X^{(2)}_{t}\dots X^{(m)}_{t})$ of $X^{(1)}_{t}$, $X^{(2)}_{t} \dots$ $X^{(m)}_{t}$. Then the expectation of the derivative with respect to time of a multivariate continuous function $\Psi(X^{(1)}_{t},X^{(2)}_{t}\dots X^{(m)}_{t})$ satisfying the conditions in Lemma 2 is of the form
$\mathbb{E}\{$ $\frac{d\Psi(X^{(1)}_{t},X^{(2)}_{t}\dots X^{(m)}_{t})}{dt}\}=<\mathcal{Y}_{X^{(1)}}\Psi,p(X^{(1)}_{t})$> (34)
@@
Proofs of Propositions 1, 3 are shown in the Appendix section of the article
By Corollary 2 and Proposition 3, it is clear that the problem of solving the derivative of the function $\Psi(X_{t})$ with respect to time $t$ is transformed, by two transformations of the $\mathcal{A}$ operator and the $\mathcal{Y}$ operator, into a problem of solving the partial differential equations for the drift and diffusion term functions of the SDEs obeyed by the independent variables in the function $\Psi(X_{t})$ with respect to their independent variables which means that solve for the three partial differential matrices $\mathcal{C}^{1}_{h,x,t}$, $\mathcal{C}^{1}_{H,x,t}$ and $\mathcal{C}^{2}_{H,x,t}$ in the $\mathcal{Y}$ operator.
Furthermore, the transformation from the $\mathcal{A}$ operator to the $\mathcal{Y}$ operator is essentially a transformation from a partial differential operator of the function $\Psi(X_{t})$ to a linear operator of the function $\Psi(X_{t})$. the $\mathcal{Y}$ operator, as a linear operator of the function $\Psi(X_{t})$, would have much better properties, both in mathematics and engineering.